# Audio Super-Resolution with Latent Bridge Models

### v2

@@ H.2.1 Audio domain
Audio waveform are continuous and high-dimensional [64], training a monolithic network to synthesize full-band signals is notoriously unstable and tends to lose high-frequency detail [16]. Recent work therefore favors cascade architectures in which each stage handles a narrower bandwidth or a simpler representation.
@@
Early systems such as Jukebox [23] map a text prompt through three VQ-VAE tiers—semantic, acoustic and waveform—before decoding; Moûsai [93] reduces this to two latent-diffusion steps. More recent models, including AudioLM, MusicLM, and MusicFlow [9, 1, 81], first generate semantic tokens, then refine them into acoustic tokens that a vocoder converts to the final audio. InspireMusic [123] pairs a transformer LM with a flow-based super-resolution decoder for long-form, high-resolution music, while YuE [121] uses an LLM-based lyric encoder followed by a lightweight vocoder, forming a two-stage lyric-to-song pipeline.
@@
On the SR side, models improve fidelity by letting each block focus on a limited frequency band: progressive up-sampling GANs (PU-GAN) [16], coarse-to-fine phase extensions such as MS-BWE [66], and multi-band diffusion that processes sub-bands in parallel [92]. The work most closely related to ours, Noise-to-Music [36], cascades two diffusion models (3.2 kHz → 16 kHz) and uses waveform blurring plus stochastic low-pass resampling to perform cascading augmentation.
@@ H.2.2 Vision domain
A similar multi-stage processing trend is evident in the vision domain. Cascade-resolution training progressively feeds the same network with increasingly finer inputs—as seen in models like PixArt-$\Sigma$[12] and SANA[116]—enabling a single model to scale seamlessly to generate high-resolution images.
@@
Modern text-to-image generation pipelines often employ cascaded architectures, where separate diffusion or autoregressive models operate sequentially at increasing spatial resolutions. Early examples include CDM [35], DALL·E 2 [83], Imagen [90], and Relay-Diffusion [104]. More recent advances, such as CogView 3 [127], Matryoshka Diffusion Model [31], and FMboost [95], incorporate hybrid approaches combining diffusion models or flow models. Pipelines like SDXL [80] enhance the image quality of Stable Diffusion [88] through explicitly incorporating multi-resolution enhancement and refinement stages.
@@
Super-resolution systems built on diffusion models also adopt cascaded frameworks. Methods such as Inf-DiT [119] and SR3 [91] perform iterative refinement, while Pixel-Space Laplacian Diffusion [2] leverage successive Laplacian pyramid-based enhancements. Together, these works confirm that allocating separate generators to coarse layout and fine detail is crucial for stable training, fast sampling, and high-fidelity ultra-resolution imagery.
@@ Appendix I Qualitative results
In this section, we present two subsections of case studies. The first subsection compares the super-resolution performance of bridge models trained in different representation spaces. The second focuses on cherry-picked samples from the demo page of A2SB [45], which provides a strong baseline for comparison.
@@ I.1 Modeling on different audio spaces
In this section, we compare three bridge models: (1) following Bridge-SR [54], we train a 10.6M-parameter network based on a large WavNet [106] architecture on our dataset; (2) following NVIDIA-NeMo [46] [^11] https://github.com/NVIDIA/NeMo, we construct a bridge model directly on the STFT spectrogram, treating each time frame as a token and modeling it with a DiT (Diffusion Transformer [4]) architecture of 0.3B, also trained on our dataset; and (3) our proposed any-to-48 kHz AudioLBM system.
@@
In this section, we present a qualitative comparison across five representative super-resolution baselines on the 8 kHz to 48 kHz task: AudioSR [62], A2SB [45], re-implemented 48 kHz version of Audit, CQTDiff [72] form from A2SB [41], and our proposed method.

### v3

@@ H.2.1 Audio domain
Audio waveform are continuous and high-dimensional [liu2023learning], training a monolithic network to synthesize full-band signals is notoriously unstable and tends to lose high-frequency detail [choprogressive]. Recent work therefore favors cascade architectures in which each stage handles a narrower bandwidth or a simpler representation.
@@
Early systems such as Jukebox [dhariwal2020jukebox] map a text prompt through three VQ-VAE tiers—semantic, acoustic and waveform—before decoding; Moûsai [schneider2024mousai] reduces this to two latent-diffusion steps. More recent models, including AudioLM, MusicLM, and MusicFlow [borsos2023audiolm, agostinelli2023musiclm, prajwal2024musicflow], first generate semantic tokens, then refine them into acoustic tokens that a vocoder converts to the final audio. InspireMusic [zhang2025inspiremusic] pairs a transformer LM with a flow-based super-resolution decoder for long-form, high-resolution music, while YuE [yuan2025yue] uses an LLM-based lyric encoder followed by a lightweight vocoder, forming a two-stage lyric-to-song pipeline.
@@
On the SR side, models improve fidelity by letting each block focus on a limited frequency band: progressive up-sampling GANs (PU-GAN) [choprogressive], coarse-to-fine phase extensions such as MS-BWE [lu2024multi], and multi-band diffusion that processes sub-bands in parallel [san2023discrete]. The work most closely related to ours, Noise-to-Music [huang2023noise2music], cascades two diffusion models (3.2 kHz → 16 kHz) and uses waveform blurring plus stochastic low-pass resampling to perform cascading augmentation.
@@ H.2.2 Vision domain
A similar multi-stage processing trend is evident in the vision domain. Cascade-resolution training progressively feeds the same network with increasingly finer inputs—as seen in models like PixArt-$\Sigma$[chen2024pixart] and SANA[xie2024sana]—enabling a single model to scale seamlessly to generate high-resolution images.
@@
Modern text-to-image generation pipelines often employ cascaded architectures, where separate diffusion or autoregressive models operate sequentially at increasing spatial resolutions. Early examples include CDM [ho2022cascaded], DALL·E 2 [ramesh2022hierarchical], Imagen [saharia2022photorealistic], and Relay-Diffusion [teng2023relay]. More recent advances, such as CogView 3 [zheng2024cogview3], Matryoshka Diffusion Model [gu2023matryoshka], and FMboost [schusterbauer2024fmboost], incorporate hybrid approaches combining diffusion models or flow models. Pipelines like SDXL [podell2023sdxl] enhance the image quality of Stable Diffusion [rombach2022high] through explicitly incorporating multi-resolution enhancement and refinement stages.
@@
Super-resolution systems built on diffusion models also adopt cascaded frameworks. Methods such as Inf-DiT [yang2024inf] and SR3 [saharia2022image] perform iterative refinement, while Pixel-Space Laplacian Diffusion [atzmon2024edify] leverage successive Laplacian pyramid-based enhancements. Together, these works confirm that allocating separate generators to coarse layout and fine detail is crucial for stable training, fast sampling, and high-fidelity ultra-resolution imagery.
@@ Appendix I Qualitative results
In this section, we present two subsections of case studies. The first subsection compares the super-resolution performance of bridge models trained in different representation spaces. The second focuses on cherry-picked samples from the demo page of A2SB [kong2025a2sb], which provides a strong baseline for comparison.
@@ I.1 Modeling on different audio spaces
In this section, we compare three bridge models: (1) following Bridge-SR [li2025bridge], we train a 10.6M-parameter network based on a large WavNet [van2016wavenet] architecture on our dataset; (2) following NVIDIA-NeMo [ku2025generative] [^11] https://github.com/NVIDIA/NeMo, we construct a bridge model directly on the STFT spectrogram, treating each time frame as a token and modeling it with a DiT (Diffusion Transformer [bao2023all]) architecture of 0.3B, also trained on our dataset; and (3) our proposed any-to-48 kHz AudioLBM system.
@@
In this section, we present a qualitative comparison across five representative super-resolution baselines on the 8 kHz to 48 kHz task: AudioSR [liu2024audiosr], A2SB [kong2025a2sb], re-implemented 48 kHz version of Audit, CQTDiff [moliner2023solving] form from A2SB [kim2024audio], and our proposed method.